9/16/16

Summary: Partly cloudy with light winds out of the NE.
                  TRH down at site 13. Tower data from Emerald channels down (refer to blog post for details) 

Actions past 24 hours:

To dos:

Sensor Status: 

T/RH:  ok, except for site 13
P: ok
2D Gill: ok
csat u,v:  ok
csat ldiag: ok
csat w, tc: ok
EC150: ok
motes: ok
Wetness: ok
radiation: ok
Tsoil: ok, bad at mote #1 (site 9)
Gsoil: ok
Qsoil: ok
Cvsoil: not responding
Vbatt: ok